Chimikin:

How do I factorize this? 1(3x - 7)-4(3x - 7)

MxxnLight:

@darkknight

Angle:

factoring means to pull out (using the distributive property) the greatest common factor For example: 4 ( x + y ) - 7 (x + y) = (x + y) ( 4 - 7)

Angle:

Another simpler example is 1x - 4x = x ( 1 - 4 ) were we take the x out of each of the terms

Angle:

@Chimikin Let me know if that helps or if you are still stuck

Chimikin:

Wow that helps thanks.

Angle:

I'm glad that helped :)

Chimikin:

:D